OpenStudy (anonymous):

Add and simplify: a+2/ 3 + a+2/ 4

OpenStudy (anonymous):

Is that: \[\frac{a+2}{3}+\frac{a+2}{4}\]

OpenStudy (anonymous):

yea :)

OpenStudy (anonymous):

im confused on how to solve it

OpenStudy (anonymous):

When you're adding fractions, you need to have the denominators be the same. Remember that anything times 1 equals itself: X*1=X Now since 1 can be written as (1/1) or (2/2) or 3/3), etc., you can rewrite the equation above like so: \[(1*\frac{a+2}{3}) + (1*\frac{a+2}{4})\]

OpenStudy (anonymous):

so what do u do after that? I'm still lost :/

OpenStudy (anonymous):

Turn each 1 into a fraction so that when you multiply it out, the bases are the same.